# Parameter estimation of range-migrating targets using OTFS signals from LEO satellites

### v1

@@ IV Rectangular shaping filters
×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k}{N}-\frac{vT}{\lambda}\right)$
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right),$ (39)
if $10\leq l^{\prime}\leq\big{\lceil}M-\frac{r\left(bT_{B}\right)M}{cT}\big{\rceil}-$ 1, and
@@
×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\left[\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k}{N}-\frac{vT}{\lambda}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]$
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l+M}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right),$ (40)
@@
For large $N/B$ and $M$, the summations in (37) only contain few dominant dominant terms; indeed, from the properties of the Dirichlet function discussed in Sec. III-A, we have $|\Psi_{k^{\prime},l^{\prime}}[k,l]|\approx 0$ if $[k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $[l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$. This result is similar to that obtained for ideal shaping filters in Remark 2. ∎
@@
×$\quad\times\Bigg{[}\sum_{l^{\prime}=0}^{M-1-l_{b,{\rm int}}-\lfloor l_{b,{\rm fra}}\rfloor}X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{b,k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}]$
@@
×$\quad\times\Bigg{[}\sum_{\bar{l}_{b}=l_{b,{\rm int}}-l}^{M-l-1-\lfloor l_{b,{\rm fra}}\rfloor}\Psi_{b,k^{\prime},l^{\prime}}[k_{\rm int}-\bar{k},l_{b,{\rm int}}-\bar{l}_{b}]$
×$\quad\times X_{\rm DD}\Big{[}[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big{]}$
@@
×$\quad\times X_{\rm DD}\Big{[}[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big{]}\Bigg{]}$
@@
×$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l_{b,{\rm int}}+l_{b,{\rm fra}}}{M}}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l-l_{b,{\rm int}}+\bar{l}_{b}}{M}}$
×$\quad\times\mathrm{e}^{\mathrm{i}2\pi\left(\frac{\bar{k}+k_{\rm fra}}{B}\right)b}A_{b,\bar{k},\bar{l}_{b}}[k,l]\mathcal{D}_{M}\!\left(\frac{\bar{l}_{b}+l_{b,{\rm fra}}}{M}\right)$
×$\quad\times X_{\rm DD}\Big{[}[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big{]},$ (41)
@@
if $-l-\lfloor l_{b,{\rm fra}}\rfloor\leq\bar{l}_{b}\leq l_{b,{\rm int}}-l-1$. If $l_{b,{\rm int}}=l_{{\rm int}}$ and $l_{b,{\rm fra}}=0$ for $b=0,\ldots,B-1$, then (41) becomes
@@
if $10\leq l\leq l_{{\rm int}}-$ 1, that is consistent with [8, Theorem 2]. ∎
@@ IV-B Approximated model
×$\quad\times\frac{1}{B}\sum_{b=0}^{B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k^{\prime}-k_{\rm int}-k_{\rm fra}}{B}\right)b}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k^{\prime}-k_{\rm int}-k_{\rm fra}}{N}\right)$
×$\quad\times\mathcal{D}_{M}\!\left(-\frac{l^{\prime}-l_{b,{\rm int}}}{M}\right)X_{\rm DD}\Big{[}[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\Big{]},$ (50)
@@
$\Theta_{k^{\prime},l^{\prime}}[k,l]=\left\{\begin{matrix}1&l^{\prime}\leq l\leq M-1,\\ \mathrm{e}^{-\mathrm{i}2\pi\frac{[k-k^{\prime}]_{N}}{N}}&0\leq l\leq l^{\prime}-1.\\ \end{matrix}\right.$ (51)
@@ V-B Extension to multiple targets
$N$ Number of symbol intervals 128
$R_{\rm r}$ Range resolution 19.5 m
$R_{\rm rr}$ Range-rate resolution 4.4 m/s
$[\bar{r}_{\rm min},\,\bar{r}_{\rm max}]$ Inspected range interval [20, 39] km
$[\bar{v}_{\rm min},\,\bar{v}_{\rm max}]$ Inspected range-rate interval [15, 16] km/s
@@ VI Numerical analysis
To assess the system performance, we consider the parameters in Table I. The maximum variation of the target range during the OTFS frame is $\bar{v}_{\rm max}NT=136$ m, corresponding to about $7$ times the range resolution $R_{\rm r}$; the initial range $d$ and the range-rate $v$ are randomly generated, while a Swerling I fluctuation model is considered for the amplitude $\alpha$. The information symbols $\{X_{\rm DD}[k,l]\}$ are independently drawn from a unit-energy $4$-QAM constellation. The entries of the observed vector $\bm{z}$ are computed according to Eqs. (11), (14), and (18). Instead, the radar receiver is designed according to the echo model proposed in Secs. III and IV for ideal and rectangular shaping filters, respectively, that relies upon the $B$-fold stop-and-go approximation. In the implementation of the proposed two-step estimator, we set $\Sigma_{\rm r}=R_{\rm r}/100$, $\Sigma_{\rm rr}=R_{\rm rr}/100$, $\tilde{G}_{\rm r}=401$, $\tilde{G}_{\rm rr}=401$, $K=4P$, and $\epsilon=1$.
@@ VII Conclusions
×$\quad\times\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n,m^{\prime}]X_{\rm TF}[n,m^{\prime}]$
@@
×$\quad\times\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n-1,m^{\prime}]X_{\rm TF}[n-1,m^{\prime}]$
@@
×$\quad\times\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{\mathrm{i}2\pi\left(\frac{nk^{\prime}}{N}-\frac{m^{\prime}l^{\prime}}{M}\right)}$
@@
×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}\right)}e^{j2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}.$ (61)
@@
×$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}}{M}}$
×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)$
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (63)
@@
×$\quad\times\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{\mathrm{i}2\pi\left(\frac{(n-1)k^{\prime}}{N}-\frac{m^{\prime}l^{\prime}}{M}\right)}$
@@
×$\quad\times\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}\right)}e^{j2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}.$ (65)
@@
×$\quad\times\sum_{p=\left\lceil M-\frac{r\left(bT_{B}\right)}{c}\frac{M}{T}\right\rceil}^{M-1}\mathcal{D}_{M}\!\left(\frac{p-M+\frac{r\left(bT_{B}\right)}{c}\frac{M}{T}-l}{M}\right)\delta\big{[}[p-M-l^{\prime}]_{M}\big{]}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p-M}{M}}.$ (66)
@@
×$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}-M}{M}}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}$
×$\quad\times\left[\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]$
×$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}+M}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (67)

### v2

@@ IV Rectangular shaping filters
$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k}{N}-\frac{vT}{\lambda}\right)$
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right),$ (39)
if $0\leq l^{\prime}\leq\big\lceil M-\frac{r\left(bT_{B}\right)M}{cT}\big\rceil-1$, and
@@
$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\left[\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k}{N}-\frac{vT}{\lambda}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]$
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l+M}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right),$ (40)
@@
For large $N/B$ and $M$, the summations in (37) only contain few dominant dominant terms; indeed, from the properties of the Dirichlet function discussed in Sec. III-A, we have $|\Psi_{k^{\prime},l^{\prime}}[k,l]|\approx 0$ if $k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$. This result is similar to that obtained for ideal shaping filters in Remark 2. ∎
@@
$\quad\times\Bigg[\sum_{l^{\prime}=0}^{M-1-l_{b,{\rm int}}-\lfloor l_{b,{\rm fra}}\rfloor}X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{b,k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}]$
@@
$\quad\times\Bigg[\sum_{\bar{l}_{b}=l_{b,{\rm int}}-l}^{M-l-1-\lfloor l_{b,{\rm fra}}\rfloor}\Psi_{b,k^{\prime},l^{\prime}}[k_{\rm int}-\bar{k},l_{b,{\rm int}}-\bar{l}_{b}]$
$\quad\times X_{\rm DD}\Big[[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big]$
@@
$\quad\times X_{\rm DD}\Big[[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big]\Bigg]$
@@
$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l_{b,{\rm int}}+l_{b,{\rm fra}}}{M}}\mathrm{e}^{\mathrm{i}2\pi\frac{k_{\rm int}+k_{\rm fra}}{N}\frac{l-l_{b,{\rm int}}+\bar{l}_{b}}{M}}$
$\quad\times\mathrm{e}^{\mathrm{i}2\pi\left(\frac{\bar{k}+k_{\rm fra}}{B}\right)b}A_{b,\bar{k},\bar{l}_{b}}[k,l]\mathcal{D}_{M}\!\left(\frac{\bar{l}_{b}+l_{b,{\rm fra}}}{M}\right)$
$\quad\times X_{\rm DD}\Big[[k-k_{\rm int}+\bar{k}]_{N},[l-l_{b,{\rm int}}+\bar{l}_{b}]_{M}\Big],$ (41)
@@
if $l-\lfloor l_{b,{\rm fra}}\rfloor\leq\bar{l}_{b}\leq l_{b,{\rm int}}-l-1$. If $l_{b,{\rm int}}=l_{{\rm int}}$ and $l_{b,{\rm fra}}=0$ for $b=0,\ldots,B-1$, then (41) becomes
@@
if $0\leq l\leq l_{{\rm int}}-1$, that is consistent with [8, Theorem 2]. ∎
@@ IV-B Approximated model
$\quad\times\frac{1}{B}\sum_{b=0}^{B-1}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k^{\prime}-k_{\rm int}-k_{\rm fra}}{B}\right)b}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k^{\prime}-k_{\rm int}-k_{\rm fra}}{N}\right)$
$\quad\times\mathcal{D}_{M}\!\left(-\frac{l^{\prime}-l_{b,{\rm int}}}{M}\right)X_{\rm DD}\Big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\Big],$ (50)
@@
[…]
@@ V-B Extension to multiple targets
NN Number of symbol intervals 128
$R_{\rm r}$ Range resolution $\approx$ 39 m
$R_{\rm rr}$ Range-rate resolution $\approx$ 8.8 m/s
$\bar{r}_{\rm min},\,\bar{r}_{\rm max}]$ Inspected range interval [20, 39] km
$\bar{v}_{\rm min},\,\bar{v}_{\rm max}]$ Inspected range-rate interval [15, 16] km/s
@@ VI Numerical analysis
To assess the system performance, we consider the parameters in Table I. The maximum variation of the target range during the OTFS frame is $\bar{v}_{\rm max}NT\approx 136$ m, corresponding to about $3.5$ times the range resolution $R_{\rm r}$; the initial range $d$ and the range-rate $v$ are randomly generated, while a Swerling I fluctuation model is considered for the amplitude $\alpha$. The information symbols $\{X_{\rm DD}[k,l]\}$ are independently drawn from a unit-energy $4$-QAM constellation. The entries of the observed vector $\bm{z}$ are computed according to Eqs. (11), (14), and (18). Instead, the radar receiver is designed according to the echo model proposed in Secs. III and IV for ideal and rectangular shaping filters, respectively, that relies upon the $B$-fold stop-and-go approximation. In the implementation of the proposed two-step estimator, we set $\Sigma_{\rm r}=R_{\rm r}/100$, $\Sigma_{\rm rr}=R_{\rm rr}/100$, $\tilde{G}_{\rm r}=401$, $\tilde{G}_{\rm rr}=401$, $K=4P$, and $\epsilon=1$.
@@ VII Conclusions
$\quad\times\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n,m^{\prime}]X_{\rm TF}[n,m^{\prime}]$
@@
$\quad\times\sum_{m^{\prime}=0}^{M-1}H_{n,m}[n-1,m^{\prime}]X_{\rm TF}[n-1,m^{\prime}]$
@@
$\quad\times\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{\mathrm{i}2\pi\left(\frac{nk^{\prime}}{N}-\frac{m^{\prime}l^{\prime}}{M}\right)}$
@@
$\quad\times\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}\right)}e^{j2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}.$ (61)
@@
$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}}{M}}$
$\quad\times\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)$
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (63)
@@
$\quad\times\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}[k^{\prime},l^{\prime}]\mathrm{e}^{\mathrm{i}2\pi\left(\frac{(n-1)k^{\prime}}{N}-\frac{m^{\prime}l^{\prime}}{M}\right)}$
@@
$\quad\times\mathrm{e}^{-\mathrm{i}2\pi n\left(\frac{k-k^{\prime}}{N}\right)}e^{j2\pi\left(\frac{ml-m^{\prime}l^{\prime}}{M}\right)}.$ (65)
@@
$\qquad\times\sum_{p=\left\lceil M-\frac{r\left(bT_{B}\right)}{c}\frac{M}{T}\right\rceil}^{M-1}\mathcal{D}_{M}\!\left(\frac{p-M+\frac{r\left(bT_{B}\right)}{c}\frac{M}{T}-l}{M}\right)\delta\big[[p-M-l^{\prime}]_{M}\big]\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{p-M}{M}}.$ (66)
@@
$\quad\times\mathrm{e}^{\mathrm{i}2\pi\frac{v}{\lambda}\frac{r\left(bT_{B}\right)}{c}}\mathrm{e}^{\mathrm{i}2\pi\frac{vT}{\lambda}\frac{l^{\prime}-M}{M}}\mathrm{e}^{-\mathrm{i}2\pi\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)\frac{bN}{B}}$
$\quad\times\left[\mathcal{D}_{\frac{N}{B}}\!\left(\frac{k-k^{\prime}}{N}-\frac{vT}{\lambda}\right)-\mathbbm{1}_{\{b=0\}}\frac{B}{N}\right]$
$\quad\times\mathcal{D}_{M}\!\left(-\left(\frac{l-l^{\prime}+M}{M}-\frac{r\left(bT_{B}\right)\Delta}{c}\right)\right).$ (67)